# Squeezing Quantum States in Three-Dimensional Twisted Crystals –ِ

### v1

@@
In a three-dimensional twisted crystal each layer sees a potential that is periodic on an emergent moir superlattice scale [andrei_marvels_2021]. In this situation the single saddle point (SP) is replaced by an energy landscape containing a translationally ordered array of compensating SP’s located at sublattice positions repeated on a Bravais superlattice (Fig. 4(a,b)). In the strong field limit, the Coriolis deflection confines states in orbits near the potential extrema. Unstable trajectories that propagate along paths connecting SP’s can escape by propagating along two-dimensional periodic networks [fertig_transmission_1987, chalker_percolation_1988, pal_emergent_2019, de_beule_network_2021]. Importantly, although the crystal potential is invariant under the network lattice translations , the coherent states accumulate path-dependent phases when displaced by ’s making them noncommutative [zak_magnetic_1964, haldane_origin_2018]. Thus the superlattice displacements are generally not identified with the set of magnetic translations [zak_magnetic_1964] that define an effective magnetic unit cell with area which is continuously tunable. The noncommutative structure of the elementary translations provides a local measure for topological order in a band even in the absence of long range crystalline order [bianco_mapping_2011, dereli_bloch_2021]. Here the effective magnetic flux penetrating a single unit cell of the moir superlattice changes as a function of the conserved twist momentum and produces quantum interference in a Hofstadter spectrum whose gap structure depends on the connectivity of the network. Two examples are given in Fig. 4 for models with square and triangular Bravais lattices. In either case the Hofstadter spectra depend on whether the intra-saddle point amplitudes (double well tunneling between a pair of equilibrium states bound near a single SP [novaes_generalized_2003]) or intra-saddle point amplitude [fertig_transmission_1987, chalker_percolation_1988] (paths connecting neighboring SP’s) are dominant. In the latter case the spectrum shows the signature of well developed resonances from weakly coupled intra-cell orbital loops linked by the effective flux. Either limit in Fig. 4 should be physically realizable, being controlled by the relative strengths of the moir crystal field, the kinetic energy penalty for confining states within a moir cell and the strength of the regularizing (pinning) potential. In a typical case, for sufficiently small rotation angle and weak pinning potential, the resonant patterns in Fig. 4(d) and 4(f) are generic. They describe a reconstruction of the low energy degrees of freedom for a nearly free particle in a twisted crystal potential that produces orbital degeneracies at the sublattice nodes of an emergent moir superlattice. Note that the edge state structure for the single SP problem (Fig. 2) reconstructs in the super moir problem so that boundary modes appear within the bulk Hofstadter gaps (Figs. S5 and S6 in the Supplementary Information).
Discussion
@@
The twist-derived gauge coupling in this model is time-reversal symmetric and by itself does not produce a Hall response: opposing contributions to response functions from twist sectors cancel. This can be avoided by biasing the populations of the orbital states with an applied magnetic field, by exciting with circularly polarized light [ji_opto-twistronic_2023] or by spontaneous symmetry breaking from the projected interactions between electrons residing in these narrow bands. Many of the features found here in three dimensions have striking counterparts in twisted two-dimensional few-layer van der Waals materials [andrei_marvels_2021]. Theory suggests and experiments observe bands of low energy degrees of freedom, possibly with topological character, spectrally isolated from remote manifolds of higher energy dispersive states. The conditions under which these low energy bands can develop coherence with the analytic structure of lowest Landau level states has been addressed in recent studies [tarnopolsky_origin_2019, ledwith_vortexability_2023]. A promising approach is to now relate the ideas developed here in three dimensions to the phenomenology of quasi two-dimensional ultrathin materials.
Author Contributions
E.J.M. and V.T.P. conceived the project and contributed to all phases of the analysis. S.A. contributed to the design and interpretion of models for the coherent state spectrum. K.K. carried out analysis of the the Bogoliubov spectra. R.J.S. contributed to symmetry analysis of the model. C.D.B. and M.M.A.Z. analyzed the Hofstadter network. All authors contributed to writing and editing the manuscript.
Acknowledgments
@@
Supplemental Material
Contents
S1 A Tale of Two Reference Frames
S1.1 Coordinate Systems
In a three-dimensional twisted crystal, there are two equivalent frames of reference using which to describe the dynamics. The first is the usual laboratory frame. This is the frame adopted in the main text because it is familiar and intuitive. It takes the perspective of an external observer. In this frame, we can either affix a rectangular coordinate system or a cylindrical coordinate system It seems only sensible to place the origin at the twist axis. The second is the crystal frame, also referred to as the twisted frame, in which the coordinates co-rotate with the crystal. This frame adopts the point of view of the crystal, not the external observer; therefore, it has the advantage that it naturally encodes the helical symmetry of the crystal, and consequently, allows certain crystal potentials to be written in a simplified manner. In this frame, we can also affix a helical coordinate system which is defined relative to the laboratory coordinates as follows [vachagina2014use]:
where is the pitch of the twist (and has dimension of inverse length). This frame is chosen such that a fixed point winds in counterclockwise fashion as increases, as shown in Fig. S1. In the laboratory frame, the Laplacian operator is simply
@@
S1.2 Dual Representations of the Free-Particle Hamiltonian
In this work, we work with Schrodinger particles with quadratic dispersion. In any frame, the free-particle Hamiltonian is just where is the mass of the particle. In the laboratory frame, the momentum operator is
@@
where we have defined new momentum operators in the twisted frame analogously to those defined in the laboratory frame. It is apparent that commutes with the Hamiltonian, so we can project into a sector of definite eigenvalue of This allows us to replace with a -number: To bring Eq. (S5) into a form similar to Eq. (S6), we define a screw operator in the laboratory frame:
In terms of the screw operator, we have Using the canomical commutation relations we find that
@@
S1.3 Plane Waves versus Screw States
In the twisted frame, we use to label eigenvalues of the operator The associated eigenfunctions are simply plane waves These functions are separable in the sense that they depend only on In the laboratory frame, we use to label eigenvalues of the operator (actually, and are the same operator just expressed in two different frames). In polar coordinates, and this operator is just
@@
where is an integer labeling the angular momentum. Rewritten in the twisted frame where these eigenvalues have the nice form (notice that this is also an eigenfunction of ). We notice that the screw states differ from the plane waves only in that they encode a helical twist explicitly. Because of that, these states depend on both and in both frames, and consequently, they depend also on angular momentum. In other words, eigenstates of couple the in-plane and out-of-plane degrees of freedom (unless the angular momentum is zero).
S2 Landau-Level Representation of the Free-Particle Hamiltonian
S2.1
In this section, we derive the Landau-level representation of the Hamiltonian. By completing the square, we find
The first two terms remind us of canonical momentum being boosted by a gauge vector field. If we define a magnetic vector potential with an associated magnetic field we find that where is the elementary charge unit. Therefore, we observe that the screw eigenvalue (or the momentum in the -direction) plays the role of a magnetic field that couples to the planar motion in the - plane. For , this magnetic field vanishes because of a decoupling between the out-of-plane and in-plane degrees of freedom. The term is a centrifugal potential that enters the Hamiltonian importantly with a negative sign irrespective of the sign of . The term is an overall shift in energy depending on the screw eigenvalue (or the vertical momentum). Ordinarily, shifts in the spectrum can be disregarded, but we caution against that temptation in this case since this is a relative shift that depends on and not an overall shift of the whole Hamiltonian. The final term measures -projected angular momentum of the wavefunction. For later convenience, we define the cyclotron frequency (this quantity is the same as in the main text, where stands for cyclotron there; here, we instead emphasize the dependence on for later convenience) and magnetic length (this quantity is the same as in the main text; here, we instead emphasize the dependence on for later convenience) in analogy to the Landau-level problem:
We now introduce the left-handed and right-handed operators mentioned in the main text:
The operators are transformed into the operators under exchanging Consequently, these operators degenerate when More seriously, the and operators completely disappear from these expressions. Therefore, the case must always be handled carefully separately. We will return to this at the end of the section. For now, ignoring these operators satisfy the following commutation relations:
Clearly, these four operators are Hermitian. We note that the commutation change sign based on Factors of are omitted in the main text for succinctness, but they need to be considered to obtain the correct algebra. Here, we carefully keep track of all the factors of We are now ready to define ladder operators:
We emphasize that these and operators are defined in a definite sector. So a different gives a different set of operators. If one insists, one can emphasize this dependence by including subscripts: and ; we leave them off to avoid cluttered notion. The factors of have been inserted specifically to ensure the usual ladder commutation relations:
The angular momentum operator in the ladder representation is
Compared to the operator defined in the main text, there is a factor of in the formula here. The harmonic potential in the ladder representation is
The momentum terms in the ladder representation simplify to
We can simplify by noting
Putting the pieces together, we find
Defining the total mode operator we can simplify the Hamiltonian even further:
It is worth pointing out that since and the whole Hamiltonian satisfies Therefore, eigenstates of can be labelled by angular momentum, which we will demonstrate explicitly later.
S2.2
For completeness, we now handle the case. We always assume for otherwise there is no twist. The Hamiltonian now is just As mentioned previously, the operators defined before degenerate and the ladder operators and no longer couple to the coordinates and So this construction does not work for the case. However, a very similar construction does work. Essentially, we define a new operator analogous to with replaced by just and a new operator analogous to also with replaced by just (this replacement is permissible since has units of )
@@
We note that Eq. (S27) is identical to Eq. (S24) with and the appropriate replacements of operators and constants. This is a nice consistency check. However, it is worth pointing out that while Eq. (S24) depends on the sign of Eq. (S27) does not depend on the sign of at all since all the factors appear as This is because when there is no propagation along the vertical direction, so the corresponding eigenstates cannot distinguish between a left-handed twist and a right-handed twist. With that important distinction made, one can now solve Eq. (S27) using the same techniques used for Eq. (S24). This completes the construction of the Landau-level representation of the free-particle Hamiltonian in all relevant scenarios.
S3 Diagonalization of the (Almost) Free-Particle Hamiltonian
S3.1 Special-Function Approach
There are different ways to diagonalize Eq. (S12). Here, we take the most elementary approach using special functions. It is simplest to work in polar coordinates where the Hamiltonian takes the form
where and One can choose to be in either the twisted frame or laboratory frame according to the discussion in Sec. S1.2. The permissible eigenfunctions and energies are
where and are the vertical and lateral sizes of the box regularization employed to normalize the eigenstates, and is a Bessel function of the first kind. It is important to note that these functions are only marginally normalizable (there are other solutions but those are not even marginally normalizable). This makes sense since we are dealing with a free particle without any confining potential. These solutions are known as Bessel beams in the context of electron vortex states [schattschneider2011theory]. The Bessel-function basis is qualitatively very different from the basis states of the operators defined in Sec. S2. While the former is extended, the latter is exponentially well-localized. To see this, we observe that basis states in Sec. S2 are:
@@
A technical clarification is in order. The appearance of does not originate directly from the algebra above. We have inserted this by hand to ensure that the resulting state remains an eigenstate of with the same eigenvalue with which we began. This is necessary because the and operators are only defined within a sector. Consequently, it cannot map states outside of that sector. If we insist that the extra exponential factors not be added ad hoc, we can modify the definition of the and operators to explicitly account for them. To do this, we note that adds unit of angular momentum. So, it should become, in the real space representation, On the other hand, depletes unit of angular momentum. So, it should become, in the real space representation, These phases clearly preserve the commutation relations, and do not change the form of the Hamiltonian. The only effect comes when writing down the wavefunctions. These phases ensure that states at with remain inside the sector when acted upon repeatedly by and We shall leave these phases off in the primary discussion, multiplying them in only when necessary. In the twisted frame, when working with and not these phases are never necessary because is completely decoupled from the in-plane degrees of freedom. This technical remark applies equally well to the case of . Now, Eq. (S34) can be simplified by the use of associated Laguerre polynomials
We thus find [robinett2006quantum]
@@
where and are the total mode number and angular momentum defined similar to before. The eigenstates are the same Gauss-Laguerre functions as in Eq. (S36) with replaced by respectively. This approach also makes it apparent that no special treatment is required for the case.
S3.2 Bogoliubov Transformation
A standard approach to solve quadratic bosonic Hamiltonians is using Bogoliubov transformations [colpa1978diagonalization, van1980note, Gurarie2003Bosonic, Complex2005Rossignoli]. Here, we use such a transformation to show that we can get exactly Eq. (S41), confirming the correctness of the result. The quadratic part of the kinetic energy can be rewritten as follows:
We note that this Hamiltonian has a block-diagonal structure because there are no terms of the sort and . A bosonic Bogoliubov transformation is defined by properties of a symplectic group. Williamson’s Theorem states that for any real positive-definite matrix, one can find a symplectic transformation such that the symplectic eigenvalues are positive-definite [nicacio2021williamson]. Unfortunately, the matrix is not positive-definite. Therefore, we need to regularize it by replacing the off-diagonal elements by This regularization turns out to be equivalent to imposing a confining potential, as we will show. We now implement the Bogoliubov transformation [Gurarie2003Bosonic]:
where we assume that and are real numbers. Next, we impose commutation relations on and , which are satisfied if we choose
The last condition can be satisfied by choosing and where is a real parameter. In terms of these new variables, we have
This matrix can be made diagonal by choosing
So is the Hermitian adjoint of and is the Hermitian adjoint of as desired. The quadratic Hamiltonian is now diagonal
Next, we need the term:
Therefore, in terms of these Bogoliubov operators, we have the following Hamiltonian
The eigenstates and energies are given by
Here, is an eigenstate of the number operator and likewise for An alternative way to label states is by their total quanta and angular momentum [footnote1]:
In either case, the eigenstates are labeled by three quantum numbers or because we are in three dimensions. This is true no matter the regularization employed because a Bogoliubov transformation must preserve mode number. Our system is a two-mode bosonic before and after performing a Bogoliubov transformation. The spectrum in Eq. (S53) is almost exactly the spectrum in Eq. (S41). To make them agree exactly, all we have to do is relate to the confining potential. In terms of the original ladder operators, the Hamiltonian with the confining potential becomes
where is defined differently from before (hence the prime). From this, we read off
and the prefactor in the energy becomes exactly as before. To calculate the wavefunctions, we compute
By noting that and we can simplify significantly
This shows that apart from factors of the and operators are the same as the and operators. It is worth pointing out that unlike in the definitions of and where the appearance of is crucial to ensure the commutation relations (because there, and are multiplied by ), the existence of here is a mere gauge choice since the operators are defined primarily with respect to the confining momentum (whose sign is unchanged) and not (whose sign does change). Therefore, and are formally equivalent up to an unimportant gauge choice, confirming that our diagonalization is correct.
S3.3 Energy Spectrum
@@
To avoid too many variables, we continue to use and to mean both dimensionless and dimensionful quantities. It should be clear from context which definition is employed in a particular instance. Recalling that and where and can be any non-negative integer, we find that must be an integer. This implies that i.e. for a given must be at least as big as Furthermore, and must be of the same parity since the sum and difference have to be divisible by Therefore, where For a fixed the range of is bounded: The bounds are saturated when we have only Bogoliubov quasiparticles of one type. Some representative spectra are shown in Fig. S2. From there, we first observe that if all values of are included, the spectra are dense in energy. This must be the case since we are plotting a four-dimensional data set projected onto a two-dimensional surface. In other words, because the energy is characterized by three quantum numbers, we know that the spectrum is dense when plotted against just one of them (in this case, ). For a fixed the minimum energy For a fixed and the band dispersion is quadratic in with vertex centered at For a fixed the spectrum at different values of are just simply shifted upward in energy by If we are interested in an energy window then roughly only modes with contribute to the physics. It makes sense that the weaker the confinement potential, the more modes must be included in the energy window.
S4 Effect of Saddle-Point Potential Energy

### v2

@@ VI Hofstadter Network
In a three-dimensional twisted crystal each layer sees a potential that is periodic on an emergent moir superlattice scale [5]. In this situation the single saddle point (SP) is replaced by periodic energy landscape containing a translationally ordered array of compensating SP’s located at sublattice positions on a Bravais superlattice (Fig. 4(a,b)). We study this in a model where an intralayer scalar potential on the moir length scale is induced by the mismatch of translation vectors on the nearest layer neighbor layers. In the strong field limit, the Coriolis deflection in such a scalar potential confines states in closed orbits near the potential extrema. Unstable trajectories that propagate along paths that connect SP’s can escape by propagating along the links of two-dimensional periodic networks [24, 25, 26, 27]. Importantly, although the crystal potential is invariant under the network lattice translations , the coherent states accumulate path-dependent phases when displaced by ’s making them noncommutative [28, 29]. Thus the superlattice displacements are generally not identified with the set of magnetic translations [28] that define an effective magnetic unit cell with area which is continuously tunable by the twist eigenvalue. This noncommutative structure of the elementary translations provides a local measure for topological dynamics within in a projected Hilbert space even in the absence of long range crystalline order [30, 31]. Note that here the effective magnetic flux that links a single unit cell of the moir superlattice changes continuously as a function of the conserved twist momentum and this produces quantum interference in a Hofstadter spectrum whose gap structure depends on the connectivity of the network. Two examples are displayed in Fig. 4 for models constructed on square and triangular Bravais lattices. In either case the structure of the Hofstadter spectra depend on whether the intra-saddle point amplitudes (double well tunneling between a pair of equilibrium states bound near a single SP [32]) or inter-saddle point amplitude [24, 25] (paths connecting neighboring SP’s) are dominant. In the latter case the spectrum shows the signature of well developed resonances from weakly coupled intra-cell orbital loops that are linked by the effective flux. Either limit in Fig. 4 should be physically realizable, being controlled by the relative strengths of the moir crystal field, the kinetic energy penalty for confining states within a moir cell and the strength of the regularizing (pinning) potential. In a typical case, for sufficiently small rotation angle and weak pinning potential, the resonant patterns in Fig. 4(d) and 4(f) should be generic. They describe a reconstruction of the low energy degrees of freedom for a nearly free particle in a twisted crystal potential that produces orbital degeneracies on the sublattice nodes of an emergent moir superlattice. Note that the edge state structure for the single SP problem (Fig. 2) reconstructs in the super moir problem so that boundary modes appear within the bulk Hofstadter gaps (Fig. 4) (see Appendix E).
VII Discussion
@@
The twist-derived gauge coupling in this model is time-reversal symmetric and by itself does not produce a Hall response: opposing contributions to response functions from twist sectors cancel. This can be avoided by biasing the populations of the orbital states with an applied magnetic field, by exciting with circularly polarized light [33] or by spontaneous symmetry breaking from the projected interactions between electrons residing in these narrow bands. Many of the features found here in three dimensions have striking counterparts in twisted two-dimensional few-layer van der Waals materials [5]. Theory suggests and experiments observe bands of low energy degrees of freedom, possibly with topological character, spectrally isolated from remote manifolds of higher energy dispersive states. The conditions under which these low energy bands can develop coherence with the analytic structure of lowest Landau level states has been addressed in recent studies [34, 35]. A promising approach is to now relate the ideas developed here in three dimensions to the phenomenology of quasi two-dimensional ultrathin materials.
Acknowledgements
@@
Appendix A A Tale of Two Reference Frames
A.1 Coordinate Systems
In a three-dimensional twisted crystal, there are two equivalent frames of reference using which to describe the dynamics. The first is the usual laboratory frame. This is the frame adopted in the main text because it is familiar and intuitive. It takes the perspective of an external observer. In this frame, we can either affix a rectangular coordinate system or a cylindrical coordinate system It seems only sensible to place the origin at the twist axis. The second is the crystal frame, also referred to as the twisted frame, in which the coordinates co-rotate with the crystal. This frame adopts the point of view of the crystal instead of that of the external observer; therefore, it naturally encodes the helical symmetry of the crystal, and consequently, allows certain crystal potentials to be written in a simplified manner. In this frame, we can also affix a helical coordinate system which is defined relative to the laboratory coordinates as follows [36]:
where is the pitch of the twist (and has dimension of inverse length). This frame is chosen such that a fixed point winds in counterclockwise fashion as increases, as shown in Fig. 5. In the laboratory frame, the Laplacian operator is simply
@@ A.1 Coordinate Systems
A.2 Dual Representations of the Free-Particle Hamiltonian
In this work, we consider only Schrodinger particles with quadratic dispersion. In any frame, the free-particle Hamiltonian is just where is the mass of the particle. In the laboratory frame, the momentum operator is
@@ A.2 Dual Representations of the Free-Particle Hamiltonian
where we have defined new momentum operators in the twisted frame analogously to those defined in the laboratory frame. It is apparent that commutes with the Hamiltonian, so we can project into a sector of definite eigenvalue of This allows us to replace with a -number: To bring Eq. (36) into a form similar to Eq. (37), we define a screw operator in the laboratory frame:
In terms of the screw operator, we have Using the canonical commutation relations we find that
@@
A.3 Plane Waves versus Screw States
In the twisted frame, we use to label eigenvalues of the operator The associated eigenfunctions are simply plane waves In the laboratory frame, we use to label eigenvalues of the operator Actually, and are the same operator just expressed in two different frames; the use of and here is to emphasize which frame we are using. In polar coordinates, and this operator is just
@@ A.3 Plane Waves versus Screw States
where is an integer labeling the angular momentum. Rewritten in the twisted frame where these eigenvalues have the nice form (notice that this is also an eigenfunction of with eigenvalue ). We notice that the screw states differ from the plane waves only in that they encode a helical twist explicitly. Because of that, in the laboratory frame, these states depend on both and , and consequently, they depend also on angular momentum. In other words, eigenstates of couple the in-plane and out-of-plane degrees of freedom (unless the angular momentum is zero). Said more simply, in the twisted frame, the eigenvalue of is just the coefficient in front of in the exponential, i.e. the wavenumber; in the laboratory frame, the eigenvalue of is not just the coefficient in front of in the exponential but the eigenvalue also depends on the angular momentum .
Appendix B Landau-Level Representation of the Free-Particle Hamiltonian
In Sec. II of the main text, we have presented the Landau-level representation of the Hamiltonian when . For completeness, we now handle the case. We always assume for otherwise there is no twist. The Hamiltonian now is just As mentioned previously, the operators defined before degenerate and the ladder operators and no longer couple to the coordinates and So this construction does not work for the case. However, a very similar construction does work. Essentially, we define a new operator analogous to with replaced by just and a new operator analogous to also with replaced by just (this replacement is permissible since has units of )
@@ Appendix B Landau-Level Representation of the Free-Particle Hamiltonian
We note that Eq. (45) is identical to Eq. (10) with and the appropriate replacements of operators and constants. This is a nice consistency check. However, it is worth pointing out that while Eq. (10) depends on the sign of Eq. (45) does not depend on the sign of at all since all the factors appear as This is because when there is no propagation along the vertical direction, so the corresponding eigenstates cannot distinguish between a left-handed twist and a right-handed twist. With that important distinction made, one can now solve Eq. (45) using the same techniques used for Eq. (10). This completes the construction of the Landau-level representation of the free-particle Hamiltonian in all relevant scenarios.
Appendix C Diagonalization of the (Almost) Free-Particle Hamiltonian
C.1 Special-Function Approach
There are different ways to diagonalize Eq. (1). Here, we take the most elementary approach using special functions. It is simplest to work in polar coordinates where the Hamiltonian takes the form
where and One can choose to be in either the twisted frame or laboratory frame according to the discussion in Sec. A.2. The permissible eigenfunctions and energies are
where and are the vertical and lateral sizes of the box regularization employed to normalize the eigenstates, and is a Bessel function of the first kind. It is important to note that these functions are only marginally normalizable (there are other solutions but those are not even marginally normalizable). This makes sense since we are dealing with a free particle without any confining potential. These solutions are known as Bessel beams in the context of electron vortex states [37]. The Bessel-function basis is qualitatively very different from the basis states of the operators defined in Sec. II. While the former is extended, the latter is exponentially well-localized. To see this, we observe that basis states in Sec. II are:
@@ C.1 Special-Function Approach
A technical clarification is in order. The appearance of does not originate directly from the algebra above. We have inserted this by hand to ensure that the resulting state remains an eigenstate of with the same eigenvalue with which we began. This is necessary because the and operators are only defined within a sector. Consequently, it cannot map states outside of that sector. If we insist that the extra exponential factors not be added ad hoc, we can modify the definition of the and operators to explicitly account for them. To do this, we note that adds unit of angular momentum. So, it should become, in the real space representation, On the other hand, depletes unit of angular momentum. So, it should become, in the real space representation, These phases clearly preserve the commutation relations, and do not change the form of the Hamiltonian. The only effect comes when writing down the wavefunctions. These phases ensure that states at with remain inside the sector when acted upon repeatedly by and We shall leave these phases off in the primary discussion, multiplying them in only when necessary. In the twisted frame, when working with and not these phases are never necessary because is completely decoupled from the in-plane degrees of freedom. This technical remark applies equally well to the case of . Now, Eq. (52) can be simplified by the use of associated Laguerre polynomials
We thus find [38]
@@
where and are the total mode number and angular momentum defined similar to before. The eigenstates are the same Gauss-Laguerre functions as in Eq. (54) with replaced by respectively. We note that this approach differs from the Bogoliubov transformation approach used in the main text in that no special treatment is required for the case. However, the two are equivalent via a gauge transformation.
C.2 Energy Spectrum
@@
To avoid too many variables, we continue to use and to mean both dimensionless and dimensionful quantities. It should be clear from context which definition is employed in a particular instance. Recalling that and where and can be any non-negative integer, we find that must be an integer. This implies that i.e. for a given must be at least as big as Furthermore, and must be of the same parity since the sum and difference have to be divisible by Therefore, where For a fixed the range of is bounded: The bounds are saturated when we have only Bogoliubov quasiparticles of one type. Some representative spectra are shown in Fig. 6. From there, we first observe that if all values of are included, the spectra are dense in energy. This must be the case since we are plotting a four-dimensional data set projected onto a two-dimensional surface. In other words, because the energy is characterized by three quantum numbers, we know that the spectrum is dense when plotted against just one of them (in this case, ). For a fixed the minimum energy For a fixed and the band dispersion is quadratic in with vertex centered at For a fixed the spectrum at different values of are just simply shifted upward in energy by If we are interested in an energy window then roughly only modes with contribute to the physics. It makes sense that the weaker the confinement potential, the more modes must be included in the energy window.
Appendix D Effect of Saddle-Point Potential Energy